OpenStudy (anonymous):

sin^-1(sin(7pi/6)) ... I got 11pi/6... is that correct?

OpenStudy (jdoe0001):

hmm, what does \(\bf \Large sin^{-1}(x)\) mean?

OpenStudy (anonymous):

same as arcsin

OpenStudy (jdoe0001):

well, yes, what does arcsin stand for?

OpenStudy (anonymous):

The inverse of sin, it's between −π/2 ≤ y ≤ π/2 on the unit circle

OpenStudy (jdoe0001):

well, ok.... what does arcsin take, an angle or just a value? and what does it return? a value or just an angle?

OpenStudy (dumbcow):

pretty basic \[\sin^{-1} (\sin x) = x\]

OpenStudy (anonymous):

Okay so @jdoe0001 I think is either -30degrees or -pi/6 depending on the answer to that question. I believe it takes a value, gives and angle so it would be -30degrees? since I'm starting with the value -1/2.

OpenStudy (jdoe0001):

well, as dumbcow said \(\bf sin^{-1}(\color{blue}{\textit{some value}}) = \theta\\ sin(\theta) = \color{blue}{\textit{some value}}\\ sin^{-1}(\color{blue}{sin(\theta)}) = \theta\)

OpenStudy (dumbcow):

i will add, arcsin can have multiple solutions...so technically your answer of 11pi/6 is still correct but so is 7pi/6

OpenStudy (anonymous):

Right! +2piK right?